User Tools

Site Tools


handover:aum007

aum007

Hobart 12m

Disk VSN: FLEXBUFF | Data volume at beginning: ? GB

  • 1900UT Exp started OK. (Arwin)
  • 1941UT dbbc25/ 2204.990000,c,32,1,agc,52,58,12938,17246,0,0 . The target value is not quite attained for this frequency (Arwin).
  • 1941UT dbbcifa/1,0,agc,2,32595,48000,1. (Arwin)
  • 1941UT PPS count is 47228, different from the wiki values (Arwin).

Katherine 12m

Disk VSN: HOB+1007| Data volume at beginning: 0 GB

  • SAR Track at 243.2048 UT. Shortened Scans: 243-2044 and 243-2052, scans removed 243-2048. (Arwin)
  • 1900UT Exp started OK. (Arwin)
  • 09:06UT - Alarm has sounded for a delay drift (appears to have gone from a stable 0.32 us to 0.38. going to reset DBBC software now. (Tiege)
  • 09:11 UT - Back up and running now 0.285 us delay, first good scan will be 244-0913 (Tiege)
  • 09:35 UT - Has drifted again, did this the other day, Delay becomes unstable while transitioning from day to evening at Katherine. Going to restart the DBBC software again. (Tiege)
  • 09:39 UT - Back on schedule, only missed 0937 and 0939, 0940 first scan back (Tiege)
  • 1020 UT - Same issue again, restarted software again… mised 1020,1024 and 1027 missed (Tiege)
  • 1110 UT - Has happened again, double checked I cant power cycle the mk5. It is still not working, just going to reset software again.. (Tiege)
  • 1121 UT - Back up, missed scans 1114, 1117, 1118 and 1120 (Tiege)
  • 1137 UT - fs has crashed, likely due to dbbc resync (happened a few experiments ago). Have restarted the field system, looks like no scans have happened since the last delay drift, so scans 1114-1138 were missed inclusive (Tiege)
  • 14:33 UT ALARM: Delay through DBBC has changed. Monitor for stability and reset DBBC if it drifts. It seems it is stable now so no need to reset. (Simin)
  • 15:23 UT drifts happened again. Have to reconfigure the dbbc. Affected scans: 244-1523 - 244-1542. (Simin)
  • 16:02 UT Field system not running! try to restart pcfs. missed scan: 244-1602 - 244-1647. (Simin)
  • 17:02 UT drifts happened again. Have to reconfigure the dbbc. Affected scans: 244-1701 - 244-1704. (Simin)

Yarragadee 12m

Disk VSN: GSSC+024 (VSN not showing)| Data volume at beginning: 0 GB

  • SAR Track at 243.2142 UT. Scans removed 243-2139, 243-2140,243-2143. (Arwin)
  • 1900UT Exp started OK. (Arwin)
/home/www/auscope/opswiki/data/pages/handover/aum007.txt · Last modified: 2018/09/01 17:10 by Atifur R. Khan